mikewwe13:

Jason paid $15.50 for 3 slices of pizza and 2 burgers. Susan paid $20 for 1 slice of pizza and 4 burgers. Write a system of equation and then determine the cost of each slice of pizza and the cost of each burger. (Do Not give me the equation as your answer. ONLY the cost) Fill in the answer for the cost ONLY using the bottom boxes(Do Not include a $ in your answer) Each slide of pizza cost $ _____________________ Each burger costs $ ____________________ ________________ ________________

mikewwe13:

@Vocaloid

Vocaloid:

let p = cost of pizza and b = cost of burger "3 slices of pizza + 2 burgers cost $15.50" translates to 3p + 2b = 15.50 "1 slice of pizza + 4 burgers cost $20.00" translates to 1p + 4b = 20 solve for p and b

mikewwe13:

35

Vocaloid:

you should have two values, one for p and one for b

mikewwe13:

5.00

Vocaloid:

check your calculations again

mikewwe13:

2.50 or 1.50

Vocaloid:

when you plug your p-value and your b-value back into the equation it should still be equal for example, when you plug in 3p + 2b it should still be equal to 15.50

Vocaloid:

if you are having trouble solving the equation I am happy to help but you need to ask/say something, I can't read your mind

mikewwe13:

Each slide of pizza cost $ 15.50 Each burger costs $ 20.00 3p + 2b = 15.50 1p + 4b = 20

Vocaloid:

15.50 and 20 are not the answers 3p + 2b = 15.50 1p + 4b = 20 start by multiplying equation 1 by 2 so you can eliminate b

mikewwe13:

b=4.45 and p=2.2

Vocaloid:

good

mikewwe13:

those are the answers ?

mikewwe13:

and are you good at history ?

Vocaloid:

yes, those are the answers, no I really can't help w/ history

mikewwe13:

that's the question

Vocaloid:

I can't view it, please try to take a screenshot if you can

mikewwe13:

y

mikewwe13:

forget the y

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mikewwe13 b=4.45 and p=2.2 \(\color{#0cbb34}{\text{End of Quote}}\) ^ these are your answers

mikewwe13:

so please explain to me for the two lines

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid *** let p = cost of pizza and b = cost of burger ***

Vocaloid:

b=4.45 and p=2.2 \(\color{#0cbb34}{\text{End of Quote}}\) ^ these are your answers